OpenStudy (anonymous):

use rational exponets to simplify ^5sqrtx^15

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

x on exponent 15/5 = x on exponent 3

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

so its 3?

jhonyy9 (jhonyy9):

yes the exponent of x

OpenStudy (anonymous):

i dont know what your saying, can u draw itfor me?

jhonyy9 (jhonyy9):

so when there is x on exponent 3/2 this sign sqrt from x on exponent 3

OpenStudy (turingtest):

\[\sqrt[5]{x^{15}}=x^3\]

OpenStudy (anonymous):

thank you turning.. and johnny

jhonyy9 (jhonyy9):

you need to know that when there is exponent by fraction always the numerator sign the exponent of the term x and the denominator sign the indice of radical

jhonyy9 (jhonyy9):

hope that is understandably

OpenStudy (turingtest):

\[\sqrt[5]{x^{15}}=x^{15/5}=x^3\]because in general\[\sqrt[b]{x^a}=x^{a/b}\]